# HaSPeR: An Image Repository for Hand Shadow Puppet Recognition

### v1

@@ 4.5 Statistical Analysis
Table 1 presents the statistical properties of the HaSPeR dataset. It tabulates the proportion of samples belonging to each of the 11 classes and their corresponding training-validation splits. The ‘Dog’ class has the highest number of images (831 samples $\approx 9.96\%$), while on the contrary, the ‘Cow’ class possesses the least number of images (635 samples $\approx 7.61\%$). This suggests that overall, hand shadow puppetry of dogs is the most popular—a fact that concurs well with our anecdotal observations while searching for hand shadow puppetry performances. The proportions of samples belonging to the ‘Crab’, ‘Elephant’, and ‘Moose’ classes ($8.33\%$, $8.81\%$, and $8.59\%$ respectively) are also slightly low due to a scarcity of performance clips starring hand shadow puppets of these classes. As evident in Table 1, the image samples are quite evenly distributed across all 11 classes. Each class has $\approx 66.72\pm 0.36\%$ samples sourced from clips of professional performers and the rest $\approx 33.27\pm 0.58\%$ samples sourced from amateur clips. In totality, we end up with 4,603 samples in the training set and 3,737 samples in the validation set, thereby partitioning HaSPeR following a 55:45 ratio.
  Models Params. Performance Metrics
Vanilla w/ Classifier Block
Top-$1$ Top-$2$ Top-$3$ Top-$1$ Top-$2$ Top-$3$
                ShuffleNetV2X10 [32] 2.3M 46.29 60.71 72.97 0.6006 0.4576 0.4225 81.64 87.26 92.23 0.8405 0.823 0.7989
ViTB16 [33] 86.6M 65.9 75.06 79.82 0.7214 0.6649 0.6462 59 68.15 73.72 0.6576 0.5965 0.5792
ViTL32 [33] 306.5M 75.94 82.82 89.08 0.7998 0.7716 0.7443 77.89 85.14 90.5 0.8158 0.7891 0.7614
AlexNet [10] 61.1M 79.04 84.05 88.19 0.826 0.8015 0.7718 81.85 88.41 92.05 0.8448 0.8261 0.8022
SqueezeNet1_1 [34] 1.2M 80.67 85.57 88.94 0.83 0.8127 0.7912 80.35 85.68 89.77 0.8296 0.8135 0.7902
MobileNetV3Small [35] 2.5M 81.08 86.13 92.18 0.8529 0.8221 0.7993 81.72 87.1 90.98 0.8513 0.8258 0.8012
ConvNextLarge [36] 197.8M 81.58 87.53 91.78 0.8441 0.8208 0.7997 84.93 88.86 92.9 0.8783 0.8522 0.8384
EfficientNetB0 [37] 5.3M 81.91 85.38 90.42 0.8625 0.829 0.8012 83.54 88.22 92.96 0.8611 0.8418 0.8231
WideResNet50_2 [38] 68.9M 83.32 88.76 95.53 0.8596 0.8416 0.8158 85.65 91.06 94.32 0.8751 0.861 0.8444
MNasNet13 [39] 6.3M 83.43 86.35 88.94 0.8495 0.8388 0.8173 82.57 88.28 92.58 0.8445 0.8315 0.8094
VGG16 [11] 138.4M 83.48 87.87 91.94 0.8462 0.8376 0.8187 84.85 90.95 94.4 0.857 0.8516 0.8364
VGG19 [11] 143.7M 83.97 90.07 93.57 0.8636 0.8449 0.8307 84.02 90.09 93.25 0.8623 0.8401 0.8274
MobileNetV3Large [35] 5.5M 84.23 90.71 94.86 0.8727 0.8468 0.83 83.32 87.79 93.57 0.8682 0.842 0.8218
EfficientNetV2S [40] 21.5M 84.72 89.24 92.15 0.8667 0.8523 0.8329 85.97 90.01 92.26 0.876 0.8628 0.8482
MobileNetV2 [41] 3.5M 84.8 90.28 94.88 0.8742 0.8563 0.8373 87.23 92.32 94.48 0.8837 0.8774 0.8668
GoogLeNet [42] 6.6M 84.93 90.12 93.84 0.8615 0.8541 0.8348 83.97 88.97 92.4 0.8554 0.8448 0.8244
SwinV2B [43] 87.9M 84.98 89.96 94.11 0.8747 0.8551 0.8439 81.34 86.88 93.22 0.853 0.8227 0.8021
ConvNeXt [36] 88.6M 85.17 89.8 95.34 0.8751 0.856 0.8431 86.75 91.03 94.46 0.8803 0.8728 0.8608
ResNet18 [44] 11.7M 85.36 92.56 94.4 0.8807 0.863 0.8411 85.84 92.61 95.21 0.8793 0.8665 0.8494
SwinB [45] 87.8M 85.84 92.72 97.35 0.8789 0.8643 0.8535 85.12 91.67 95.23 0.8772 0.8564 0.843
ResNet101 [44] 44.5M 86.37 93.17 95.85 0.8837 0.8706 0.8508 86.54 93.97 97 0.8764 0.8719 0.8556
MaxViT [46] 30.9M 86.54 93.84 97.05 0.8903 0.8691 0.8642 85.2 91.54 95.18 0.8722 0.857 0.847
RegNetX32GF [47] 107.8M 86.67 91.14 95.63 0.8879 0.8735 0.857 86.99 92.13 95.21 0.8823 0.8742 0.8608
DenseNet201 [48] 20.0M 86.83 90.23 92.74 0.8853 0.8701 0.8548 88.89 93.49 95.42 0.9042 0.8953 0.8844
ResNet152 [44] 60.2M 86.86 90.42 94.4 0.8929 0.877 0.8585 86.64 91.57 94.51 0.8844 0.8715 0.8576
ResNet50 [44] 25.6M 86.91 92.74 95.47 0.8953 0.8771 0.8607 84.69 90.12 94.08 0.8717 0.8538 0.8357
ResNet34 [44] 21.8M 87.55 92.93 95.55 0.8943 0.8809 0.8699 85.79 91.73 93.79 0.8767 0.8634 0.8461
WideResNet101_2 [38] 126.9M 87.77 92.37 95.1 0.8955 0.8833 0.8684 86.27 92.45 96.14 0.8783 0.869 0.8533
DenseNet121 [48] 8.0M 87.9 92.4 94.75 0.8947 0.8844 0.8719 86.08 90.95 94.64 0.8771 0.866 0.8498
ResNext101_32X8D [49] 88.8M 88.62 93.17 95.85 0.9045 0.8927 0.8803 86.54 93.97 97 0.9073 0.896 0.8839
InceptionV3 [50] 27.2M 88.97 92.77 94.19 0.8979 0.8943 0.8828 88.3 92.29 95.34 0.8905 0.8849 0.8777
@@ 5.1.3 Classifier Network
We adopt two approaches to arrive at the final 11-dimensional layer since there are a total of 11 classes to predict from. The first approach is to directly append an 11-dimensional fully connected layer at the tail-end of the vanilla models. The second approach incorporates the classifier block portrayed in Figure 6.
@@ 5.1.4 Hyperparameters and Optimizer
We use Stochastic Gradient Descent (SGD) [53], with a learning rate $\alpha=0.001$ and momentum $\gamma=0.9$, as the optimizing method, and Cross Entropy Loss as the loss metric for all the models. To decay the learning rate, we use Step Scheduler, which decays $\alpha$ by $0.1$ every 5 epochs. Each model undergoes training for 50 epochs to ensure equitable comparison, and we empirically ascertain that 50 epochs are sufficient for the majority of the models to achieve convergence.
@@ 5.1.5 Data Augmentation and Preprocessing
In order to generate a more diverse pool of training samples, we also incorporate data transformation techniques [^10] https://pytorch.org/vision/stable/transforms.html—Random Resize, Random Perspective, Color Jitter, Random Horizontal Flip, Random Crop, Random Rotation, Gaussian Blur, and Random Affine with translation and shearing—while training the models. We choose these data augmentation techniques since the classes in HaSPeR are mostly rotationally asymmetric and incongruent. Consequently, the augmented samples aid in eliciting better generalization abilities and robustness for all the models. The input images that are fed to the models are appropriately resized a priori using Bicubic Interpolation.
@@ 5.2.1 Performance Analysis
The InceptionV3 model yielded the best performance with a top-$1$ accuracy of $88.97\%$. The vanilla version of the model also yields reasonably high Precision, Recall, and F1-score of $0.8979$, $0.8943$, and $0.8828$ respectively. In terms of top-$2$ and top-$3$ accuracy values, we observe the competence of transformer architectures like SwinB and MaxViT, with each achieving the highest top-$2$ and top-$3$ accuracy scores, respectively. Upon being equipped with the classifier block shown in Figure 6, the DenseNet201 model yields the highest top-$1$ accuracy and F1-score. In contrast, the ResNext101_32X8D model demonstrates the best performance across all the other evaluation metrics. At this recess of the performance analysis, we consider the top-$1$ accuracy metric to be the most statistically significant metric. As evident in Table 2, the ResNext101_32X8D model lags behind the InceptionV3 model when it comes to the top-$1$ accuracy value in both cases, which is why we adjudicate that the latter is the best-performing model.
@@
Figure 8: $t$-Distributed Stochastic Neighbor Embedding ($t$-SNE) feature representations of InceptionV3.
It is noteworthy to point out that MobileNetV2, with only 3.5 million parameters, managed to surpass many of the other models in terms of performance. This indicates the suitability of this image classification task for lighter, low-latency models that can be used in mobile applications and embedded devices. We create a simple prototype Android application using Flutter to test the efficacy of MobileNetV2 in classifying hand shadow puppet images from the phone’s camera feed. Figure 7 portrays the snapshots of the prototype application.
Figure 9: The juxtaposition of original image samples from the HaSPeR dataset with their corresponding GradCAM Heatmaps, LIME Visualizations, and Saliency Maps (for the best-performing vanilla InceptionV3 model).
5.2.2 Feature Space Visualization and Analysis
In order to visualize the learned feature space of InceptionV3, we resort to the dimensionality reduction technique called $t$-Distributed Stochastic Neighbor Embedding [54] since it can preserve the proximity of high-dimensional data points. For high-dimensional data residing on or proximate to a low-dimensional, non-linear manifold, it becomes imperative to preserve this proximity of the collapsed low-dimensional representations for closely resembling data points. Achieving such proximity preservation is often unattainable through linear mappings such as Principal Component Analysis (PCA), which is why we opt for the $t$-SNE dimensionality reduction approach. We can pragmatically infer from the 2D-collapsed visualizations of the high-dimensional feature representations in Figure 8, that the classes are nicely clustered and congealed with minimal overlaps and outliers. This enables the model to easily determine the decision surface in the high-dimensional feature space and perform very well on the classification task.
@@ 5.2.2 Feature Space Visualization and Analysis
5.2.3 Qualitative Analysis and Explainability
As depicted in Figure 9, we adopt a plethora of Explainable AI (xAI) techniques for the best-performing InceptionV3 model to understand its decision-making. While viewing the GradCAM (Gradient-weighted Class Activation Mapping) [55] attention heatmaps, it becomes apparent that the model puts more gravitas on the common-sense distinguishing traits. For example, in Figure 9(b), we observe the regions of the image samples predominantly influencing their respective classification scores—the wingspan and beak of a bird, the gallinaceous comb of a chicken, the horns and concave head of a cow, the appendages of a crab, the horns of a deer, the long-slanted head of a dog, the tusks of an elephant, the upright horns of a moose, the big eyes and small ears of a panther, the petite hands and head of a rabbit, as well as the shell and antennae of a snail. As human beings, we evoke these same distinguishing characteristics while classifying the images using our own visual reasoning faculties. As exemplified in Figure 9(c), for local interpretation, we use the model-agnostic technique called LIME (Local Interpretable Model-agnostic Explanations) [56]. We also demonstrate the spatial support of the top-$1$ predicted classes by generating the saliency maps [57] in Figure 9(d). These maps are rendered using a solitary back-propagation pass through the InceptionV3 model, and they accentuate the salient areas of the given image, characterized by their discriminative attributes with respect to the given class.
Figure 11: Confusion Matrix of vanilla InceptionV3.

### v3

@@ 4.5 Statistical Analysis
Table 1 presents the statistical properties of the HaSPeR dataset. It tabulates the proportion of samples belonging to each of the 15 classes and their corresponding training-validation splits. While searching for hand shadow puppetry performances, we anecdotally observe that certain classes of puppets are more popular than others. Irrespective of this fact, we make sure all 15 classes of puppets are equitably represented in our dataset, with each class having 1,000 image samples ($\approx 6.\dot{6}\%$). As evident in Table 1, the image samples are evenly distributed across all 15 classes. The proportions of professionally sourced samples belonging to the ‘Llama’ and ‘Snake’ classes ($14\%$ and $27.3\%$ respectively) are slightly low due to a scarcity of performance clips starring hand shadow puppets of these classes. For cases such as these, we supplement the classes with samples organically created by our novice shadowgraphist cohort. Each class has $\approx 47.827\pm 1.414\%$ samples from clips of professional performers, and the rest $\approx 52.172\pm 1.414\%$ samples are sourced from amateur clips. Of the 15,000 samples, approximately $76.53\%$ feature male hands, while $23.64\%$ represent female hands. In tandem, considering the target demographic of ombromanie, around $28.33\%$ of the samples in HaSPeR consist of children’s hands. In totality, we end up with 9,000 samples in the training set and 6,000 samples in the validation set, thereby partitioning HaSPeR by maintaining a 60:40 ratio.
@@ 5.1.3 Classifier Network
We adopt two approaches to arrive at the final 15-dimensional layer since there are a total of 15 classes to predict from. The first approach is to directly append a 15-dimensional fully connected layer at the tail-end of the vanilla models. The second approach incorporates the classifier block portrayed in Figure 7.
@@ 5.1.4 Hyperparameters and Optimizer
We use Stochastic Gradient Descent (SGD) [33], with a learning rate $\alpha=0.001$ and momentum $\gamma=0.9$, as the optimizing method, and Cross Entropy Loss as the loss metric for all the models. To decay the learning rate, we use Step Scheduler, which decays $\alpha$ by $0.1$ every 5 epochs. Each model undergoes training for 50 epochs to ensure equitable comparison, and we empirically ascertain that 50 epochs are sufficient for all of the models to achieve convergence.
  Models Params. Performance Metrics
Vanilla w/ Classifier Block
Top-$1$ Top-$2$ Top-$3$ Top-$1$ Top-$2$ Top-$3$
                ShuffleNetV2X10 [34] 2.3M 61.73 78.41 86.1 0.6559 0.6173 0.597 88.73 93.98 96.1 0.8995 0.8873 0.8853
ViTB16 [35] 86.6M 69.71 77.6 83.28 0.7276 0.6972 0.6969 68.88 76.65 81.36 0.7192 0.6868 0.6851
ViTL32 [35] 306.5M 85.1 91.56 94.48 0.872 0.851 0.8509 84.71 91.8 94.08 0.8632 0.8472 0.8465
AlexNet [9] 61.1M 87.01 93.61 95.46 0.884 0.8702 0.8708 88.18 92.58 94.8 0.8887 0.8818 0.8809
SqueezeNet1_1 [36] 1.2M 87.56 92.45 94.15 0.888 0.8757 0.8744 86.21 92.48 94.65 0.8754 0.8622 0.8637
MobileNetV3Small [37] 2.5M 89.48 94.31 95.76 0.9038 0.8948 0.8942 89.85 94.35 96.48 0.9082 0.8985 0.8976
SwinB [38] 87.8M 90.5 95.38 97.4 0.9128 0.905 0.9042 90.2 95.4 97.08 0.9097 0.902 0.9006
GoogLeNet [39] 6.6M 90.73 94.65 95.7 0.9105 0.9073 0.9059 92.18 95.65 96.58 0.9283 0.9218 0.9206
ResNet18 [40] 11.7M 90.91 95.28 96.6 0.9176 0.9092 0.9069 91.25 95.43 97.05 0.9229 0.9125 0.9119
MobileNetV3Large [37] 5.5M 91.2 94.48 95.98 0.9185 0.912 0.911 90.4 94.53 95.26 0.9147 0.904 0.9024
ConvNeXt [41] 88.6M 91.46 96.33 98.05 0.922 0.9147 0.914 92.55 96.36 97.96 0.9306 0.9255 0.9246
SwinV2B [42] 87.9M 91.58 96.25 97.61 0.921 0.9158 0.9151 91.48 96 97.55 0.9209 0.9148 0.9144
VGG16 [10] 138.4M 91.61 95.08 96.65 0.9248 0.9162 0.9168 91 95.21 96.45 0.9235 0.91 0.9119
MNasNet13 [43] 6.3M 91.66 95.65 97.01 0.924 0.9167 0.9149 91.45 95.86 97.26 0.9231 0.9145 0.9133
ConvNextLarge [41] 197.8M 91.88 95.9 97.7 0.9254 0.9188 0.9181 88 94.7 96.56 0.8942 0.88 0.8782
EfficientNetB0 [44] 5.3M 91.93 95.26 96.71 0.9257 0.9193 0.9178 90.4 93.75 95.1 0.9131 0.904 0.9022
MaxViT [45] 30.9M 92.01 96.5 97.81 0.9268 0.9202 0.9214 92.08 95.98 97.36 0.932 0.9208 0.9237
EfficientNetV2S [46] 21.5M 92.31 95.75 96.76 0.9375 0.9232 0.9245 94.45 97.35 98.3 0.9498 0.9445 0.9438
VGG19 [10] 143.7M 92.36 95.13 96.1 0.9354 0.9237 0.9242 91.8 95.06 96.15 0.9296 0.918 0.9187
MobileNetV2 [47] 3.5M 92.38 94.98 96.05 0.9303 0.9238 0.9233 92.31 95.38 96.91 0.9311 0.9232 0.9225
WideResNet50_2 [48] 68.9M 92.46 96.28 97.28 0.9331 0.9247 0.9235 93.35 95.73 97.15 0.9421 0.9335 0.933
ResNet50 [40] 25.6M 92.58 95.56 96.75 0.9332 0.9258 0.9252 93.08 96.48 97.2 0.9363 0.9308 0.9299
RegNetX32GF [49] 107.8M 92.86 95.71 96.93 0.9348 0.9287 0.9269 92.91 95.71 96.95 0.9366 0.9292 0.9282
DenseNet121 [50] 8.0M 92.93 95.75 96.88 0.9367 0.9293 0.9282 92.95 95.51 96.56 0.936 0.9295 0.9285
ResNext101_32X8D [51] 88.8M 93 96.41 97.23 0.9364 0.931 0.9303 94.2 96.61 97.58 0.952 0.942 0.9423
WideResNet101_2 [48] 126.9M 93.36 95.81 96.9 0.9423 0.9337 0.9332 92.73 96.35 97.63 0.9337 0.9273 0.9267
InceptionV3 [52] 27.2M 93.5 96.48 97.35 0.9401 0.935 0.9338 93.71 96.36 97.06 0.9446 0.9372 0.9371
DenseNet201 [50] 20.0M 93.56 95.78 96.73 0.945 0.9357 0.9353 94.43 97 97.61 0.9492 0.9443 0.9442
ResNet101 [40] 44.5M 93.81 96.23 97.71 0.9432 0.9382 0.9406 93.23 96.93 98.13 0.9386 0.9323 0.9321
ResNet152 [40] 60.2M 94.06 97.06 98.05 0.9447 0.9407 0.9394 93.05 96.73 97.48 0.9374 0.9305 0.9297
ResNet34 [40] 21.8M 94.97 97.23 98.23 0.9516 0.9497 0.9491 91.98 95.95 97.2 0.9266 0.9198 0.9189
@@ 5.1.5 Data Augmentation and Preprocessing
In order to generate a more diverse pool of training samples, we also incorporate data transformation techniques [^10] https://pytorch.org/vision/stable/transforms.html—Random Resize, Random Perspective, Color Jitter, Random Invert, Random Horizontal Flip, Random Crop, Random Rotation, Gaussian Blur, and Random Affine with translation and shearing—while training the models. We choose these data augmentation techniques since the classes in HaSPeR are mostly rotationally asymmetric and incongruent. Consequently, the augmented samples aid in eliciting better generalization abilities and robustness for all the models. The input images that are fed to the models are appropriately resized a priori using Bicubic Interpolation.
@@ 5.2.1 Performance Analysis
The ResNet34 model yields the best performance with a top-$1$ accuracy of $94.97\%$. The vanilla version of the model also yields the highest top-$2$ accuracy, top-$3$ accuracy, Precision, Recall, and F1-scores of $97.23\%$, $98.23\%$, $0.9516$, $0.9497$, and $0.9491$ respectively. Upon being equipped with the classifier block shown in Figure 7, the EfficientNetV2S model yields the highest top-$k$ accuracies and Recall. In contrast, the ResNext101_32X8D and DenseNet201 models demonstrate the best performance across Precision and F1-score metrics respectively. At this recess of the performance analysis, we consider the top-$1$ accuracy metric to be the most statistically significant metric. As evident in Table 2, the vanilla models listed in the penultimate row and above lag behind the ResNet34 model when it comes to the top-$1$ accuracy value (as well as the other metrics), which is why we adjudicate that ResNet34 is the best-performing model.
Figure 8: $t$-Distributed Stochastic Neighbor Embedding ($t$-SNE) feature representations of ResNet34.
5.2.2 Feature Space Visualization and Analysis
In order to visualize the learned feature space of ResNet34, we resort to the dimensionality reduction technique called $t$-Distributed Stochastic Neighbor Embedding [53] since it can preserve the proximity of high-dimensional data points. For high-dimensional data residing on or proximate to a low-dimensional, non-linear manifold, it becomes imperative to preserve this proximity of the collapsed low-dimensional representations for closely resembling data points. Achieving such proximity preservation is often unattainable through linear mappings such as Principal Component Analysis (PCA), which is why we opt for the $t$-SNE dimensionality reduction approach. We can pragmatically infer from the 2D-collapsed visualizations of the high-dimensional feature representations in Figure 8, that the classes are nicely clustered and congealed with minimal overlaps and outliers. This enables the model to easily determine the decision surface in the high-dimensional feature space and perform very well on the classification task.
5.2.3 Qualitative Analysis and Explainability
As depicted in Figure 10, we adopt a plethora of explainable AI (xAI) techniques for the best-performing ResNet34 model to understand its decision-making. While viewing the GradCAM (Gradient-weighted Class Activation Mapping) [54] attention heatmaps, it becomes apparent that the model puts more gravitas on the common-sense distinguishing traits. For example, in Figure 10(b), we observe the regions of the image samples predominantly influencing their respective classification scores—the wingspan and beak of a bird, the gallinaceous comb of a chicken, the horns and concave head of a cow, the appendages of a crab, the horns of a deer, the long-slanted head of a dog, the tusks of an elephant, the long maxilla-mandibular jaw of a horse, the long-eared and tapered head of a llama, the upright horns of a moose, the big eyes and small ears of a panther, the petite hands and head of a rabbit, the shell and antennae of a snail, the lateral hood expansion of a snake, as well as the slender neck and wing feathers of a swan. As human beings, we evoke these same distinguishing characteristics while classifying the images using our own visual reasoning faculties. As exemplified in Figure 10(c), for local interpretation, we use the model-agnostic technique called LIME (Local Interpretable Model-agnostic Explanations) [55]. The green highlights indicate regions of the image that contribute positively to the probability of the assigned label, while the red highlights signify areas that reduce this probability. We also demonstrate the spatial support of the top-$1$ predicted classes by generating the saliency maps [56] in Figure 10(d). These maps are rendered using a solitary back-propagation pass through the ResNet34 model, and they accentuate the salient areas of the given image, characterized by their discriminative attributes with respect to the given class.
@@ 5.2.3 Qualitative Analysis and Explainability
Figure 10: The juxtaposition of original image samples from the HaSPeR dataset with their corresponding GradCAM Heatmaps, LIME Visualizations, and Saliency Maps (for the best-performing vanilla ResNet34 model).
5.2.4 Practicality Analysis as a Teaching Tool
It is noteworthy to point out that MobileNetV2, with only 3.5 million parameters, managed to surpass many of the other models in terms of performance. This indicates the suitability of this image classification task for lighter, low-latency models that can be used in mobile applications and embedded devices. We create a simple prototype Android application using Flutter to test the efficacy of MobileNetV2 in classifying hand shadow puppet images from the phone’s camera feed. In order to make the prototype work as seamlessly as intended, we make sure the vicinity is well-lit, and the camera accurately captures a sharply focused silhouette. We find that the model has a memory footprint of $29$ MB and achieves an average inference time of $880$ $\mu$s on the Snapdragon 8 Gen 2 mobile chipset featured in the Samsung Galaxy S23 smartphone. Figure 9 portrays the snapshots of the prototype application. There are several other practical implementation challenges involved in this endeavor that we can identify for an educational mobile application to comport well with the target demographics and real-world settings. The system must operate with minimal computational overhead, ensure real-time responsiveness, and maintain low latency. This necessitates a pragmatic tradeoff between the FLOP count and classification accuracy, with the requisite model compression and optimization techniques. Given the potential variability in device camera capabilities, the application must have preprocessing steps including, but not limited to, denoising, adaptive contrast enhancement, and sharpening the input feed to mitigate artifacts. In recognition of the diverse motor capabilities of users, particularly younger learners and users with dexterity impairments, the application must have intelligent motion compensation to stabilize the shaky camera inputs. Adding to the desiderata is the consideration of a suitable UI/UX that is tailored for the pediatric user base, as is done in handwriting teaching apps for children [57], because an intuitive and enjoyable learning experience is of paramount importance for educational apps. This includes providing step-by-step tutorials, interactive guides with progressive difficulty, and illustrative diagrams to demonstrate the creation of hand shadow puppets. A holistically sound approach to this application development endeavor will surely potentiate the pedagogical feasibility of ombromanie.